### v2

# Permutation Learning with Only N Parameters: From SoftSort to Self-Organizing Gaussians

Abstract
Sorting and permutation learning are key concepts in optimization and machine learning, especially when organizing high-dimensional data into meaningful spatial layouts. The Gumbel-Sinkhorn method, while effective, requires parameters to determine a full permutation matrix, making it computationally expensive for large datasets. Low-rank matrix factorization approximations reduce memory requirements to (with ), but they still struggle with very large problems. SoftSort, by providing a continuous relaxation of the argsort operator, allows differentiable 1D sorting, but it faces challenges with multidimensional data and complex permutations. In this paper, we present a novel method for learning permutations using only parameters, which dramatically reduces storage costs. Our method extends SoftSort by iteratively shuffling the indices of the elements and applying a few SoftSort optimization steps per iteration. This modification significantly improves sorting quality, especially for multidimensional data and complex optimization criteria, and outperforms pure SoftSort. Our method offers improved memory efficiency and scalability compared to existing approaches, while maintaining high-quality permutation learning. Its dramatically reduced memory requirements make it particularly well-suited for large-scale optimization tasks, such as ”Self-Organizing Gaussians”, where efficient and scalable permutation learning is critical.
Index Terms:
Permutation Learning, Visual Image Exploration, Sorted Grid LayoutsI Introduction & Related Work
Learning permutations is a fundamental challenge in machine learning, computer vision, and optimization. Many real-world problems, such as ranking, assignment, and sorting, require finding an optimal arrangement of elements [16, 15, 13]. However, learning permutations efficiently is computationally challenging due to the discrete nature of permutation matrices, which prevents direct gradient-based optimization.
While our proposed permutation learning method has broad applicability across various fields, this paper focuses on its use in distance-preserving grid layout algorithms for visual image sorting and 3D scene reconstruction in 3D Gaussian Splatting [7]. Sorting images based on similarity enhances visual perception, enabling users to efficiently explore hundreds of images simultaneously. In 3D scene reconstruction, organizing each scene attribute into a sorted 2D grid increases spatial correlation, and the resulting high-dimensional maps can be compressed using standard codecs, achieving substantial reductions without sacrificing rendering quality.
I-A Permutation Learning
One of the most well-established approaches to differentiable permutation learning is the Gumbel-Sinkhorn method [11], which relaxes discrete permutations into doubly stochastic matrices and applies Sinkhorn normalization. This allows learning of approximate permutations in an end-to-end differentiable manner. However, this approach requires storing and computing an matrix with being the number of elements to be sorted, making it impractical for large-scale problems.
To reduce memory costs, low-rank factorization techniques have been proposed, such as the ”Kissing to Find a Match” method [4]. These methods approximate permutation matrices using a low-rank decomposition, requiring only parameters (with ). While this significantly reduces memory usage compared to Gumbel-Sinkhorn, it may still be too memory-intensive for very large problems.
An alternative approach is SoftSort [14], a continuous relaxation of the argsort operator that enables differentiable sorting. Unlike Gumbel-Sinkhorn, SoftSort requires only parameters, making it highly memory-efficient. However, its inherent limitation to 1D sorting restricts its use in complex permutation tasks like assignment problems or distance-preserving grid layouts.
In this paper, we extend SoftSort to learn permutations with only parameters while overcoming its quality limitations and enabling multi-dimensional sorting.
I-B Heuristic Distance-Preserving Grid Layout Algorithms
Distance-Preserving Grid Layouts optimize the assignment of objects or images to grid positions by aligning spatial proximity with the similarity relationships of their feature vectors. Various algorithms have been proposed for arranging vectors on a 2D grid, most of which rely on heuristic methods that offer efficient performance. However, recent research [2] has demonstrated that gradient-based learning techniques can produce sorted grid layouts with superior sorting quality compared to traditional approaches. In the following sections, we present a brief overview of the most commonly used non-learning methods.
A Self-Organizing Map (SOM) [8, 9] uses a grid of map vectors with the same dimensionality as the input vectors. It assigns these vectors to the most similar grid position and iteratively updates the map vectors based on their neighborhood.
A Self-Sorting Map (SSM) [17, 18] initializes grid cells with input vectors and employs a hierarchical swapping process. This method compares vector similarity with the average of their grid neighborhood, considering all possible swaps within a set of four cells.
Linear Assignment Sorting (LAS) [3] merges the concepts of SOM (using a continuously filtered map) and SSM (performing cell swaps), extending the idea to optimally swap all vectors simultaneously. Fast Linear Assignment Sorting (FLAS) improves runtime efficiency by iterative swapping subsets, achieving sorting quality close to LAS.
Dimensionality reduction methods like t-SNE [19] and UMAP [10] can project high-dimensional vectors onto a 2D plane before grid arrangement. Linear assignment solvers, such as the Jonker–Volgenant algorithm [6], then map these positions to optimal grid locations. Fast placement strategies are discussed in [5].
I-C Contributions
In [2], we introduced the first gradient-based grid sorting method using Gumbel-Sinkhorn. It ensures a valid permutation matrix while optimizing the grid arrangement for vector similarity, achieving superior sorting quality. However, its memory demand limits scalability for large datasets.
To address this challenge, we introduce ShuffleSoftSort, an extension of SoftSort designed to improve sorting performance with limited memory requierements.
ShuffleSoftSort requires only parameters. By using an iterated shuffling of the sorting indices, it iteratively reorders elements along different pathes, enabling large-scale swaps not possible with SoftSort.
We show that ShuffleSoftSort achieves better performance to low-rank methods such as Kissing to Find a Match at reduced memory requirements.
By drastically reducing memory requirements compared to Gumbel-Sinkhorn while maintaining strong sorting performance, ShuffleSoftSort enables scalable permutation learning for large-scale optimization problems such as Self-organizing Gaussians or data visualization.
II Proposed Method
Figure 2 illustrates the principle of permutation learning with the goal of estimating a permutation matrix that sorts the input vectors into . Since is not differentiable, a continuous version is optimized by adjusting the weights based on a given loss function. The loss function enforces that , a binarized version of , constitutes a valid permutation while meeting the optimization requirements.
The following section discusses various permutation approximation approaches, beginning with existing methods and concluding with the proposed ShuffleSoftSort technique.
Gumbel-Sinkhorn uses iterative normalization of rows and columns of a matrix formed from the weight values, to ultimately produce a doubly stochastic matrix. The Gumbel-Sinkhorn operator introduces Gumbel noise to this process, allowing for the approximation of a permutation matrix through a differentiable relaxation. A scaling factor controls the sharpness of the approximation, with smaller values of resulting in a closer approximation to a permutation matrix. Gumbel-Sinkhorn can generate high-quality permutations but has quadratic memory consumption.
To mitigate the high memory requirements of large permutation matrices, Droge et al. propose in their paper ”Kissing to Find a Match” [4] a method based on low-rank matrix factorization. This approach involves two row-normalized matrices and of size , where . The permutation matrix is then approximated by , followed by scaling and row-wise softmax normalization.
Our new approach is based on SoftSort, which was originally developed as a continuous relaxation for the argsort operator, defined as
with being the L1 distance matrix between the sorted and the unsorted elements of the vector . Row-wise softmax increases matrix elements at positions where elements coincide, while suppressing the others. With decreasing the matrix converges to the permutation matrix.
SoftSort produces valid permutations without iterated normalization, but is poorly suited for high-dimensional data due to its inherent one-dimensional sorting.
Figure 1 (left) shows a 2D sorting of RGB colors with SoftSort. Since the order of the elements in SoftSort is represented by a one-dimensional weight tensor, each change of a weight during learning leads to a shift of the element to be sorted along this one-dimensional order.
Figure 3 illustrates the SoftSort problem using a 1D toy example. If the yellow and magenta colors were swapped, the color ordering based on minimal neighbor distances would be perfect. However, reaching the correct positions would require passing through intermediate colors, causing gradient-based optimization to fail, as the quality initially degrades. A similar issue occurs in Figure 1, where swapping elements separated by multiple rows encounters the same problem.
The following table summarizes the properties of the discussed permutation approximation methods, along with the characteristics of our proposed ShuffleSoftSort scheme. Stability here means how certain it is to obtain a valid permutation matrix without duplicates.
Gumbel- Sinkhorn | Kissing | SoftSort | Ours | |
Number of Parameters K | ||||
Non-iterative normalization | no | yes | yes | yes |
Quality | ++ | + | – | ++ |
Stability | + | o | ++ | ++ |
Our newly proposed approach, ShuffleSoftSort, retains the advantages of SoftSort, such as low memory usage and the ability to easily achieve valid permutations, while improving the quality of the permutations compared to SoftSort. Since SoftSort can only perform one-dimensional sorting, it struggles with problems like those shown in Figures 1 or 3. To solve this problem, we circumvent the one-dimensional constraint without altering the fundamental principle.
Our solution is to iteratively reorganize/shuffle the indices of the data to be sorted and then apply SoftSort using the updated one-dimensional indices. Rearranging the 1D order allows elements to be repositioned that would not have been possible in the original order.
Figure 4 illustrates the core concept of ShuffleSoftSort, highlighting how the alternating shuffled index organization improves sorting flexibility. Algorithm 1 describes the ShuffleSoftSort procedure for sorting the elements of a vector with shape .
The loss function is computed on the reverse-shuffled elements. Simply shuffling and repeatedly applying SoftSort is ineffective, as each iteration would undo the previous sorting. To address this, the weights in each SoftSort iteration must be set in a manner that initially preserves the previous order. This can be achieved by initializing the weights in a linear ascending order, combined with a small initial temperature parameter . During training, is gradually reduced to gradually reduce the influence of more distant neighbor elements. The final achievable sorting quality increases with the number of iterations .
This new sorting strategy extends to other permutation learning tasks, with the key idea of continuously modifying the index order to allow for greater flexibility in sorting. For a memory-efficient implementation, it is crucial to compute the permutation matrix and the loss elements in a row-wise manner, as the complete storage would require elements.
For grid-based sorting with ShuffleSoftSort, we use a loss function that can be efficiently computed in a separable manner. The loss builds on [2], incorporating the neighborhood loss (the normalized average distance of neighboring grid vectors in horizontal and vertical directions). However, it avoids the computationally and memory-intensive distance matrix loss. Our used loss function is:
We obtained the best results by combining two loss components to generate valid permutation matrices. The first regularization term, the stochastic constraint loss , ensures that converges to a doubly stochastic matrix by penalizing deviations from 1 in the column sums of the matrix:
The second term, , is the standard deviation loss, which encourages to converge to a permutation matrix by ensuring that the standard deviation of the original and sorted vectors remains the same:
The regularization parameters and were set to 1 and 2 respectively. SoftSort typically produces a valid permutation matrix. In very rare cases, where the columns of the permutation matrix contain duplicates, the SoftSort iterations are extended until a valid permutation is achieved.
III Experimental Evaluation
In this section, we compare Gumbel-Sinkhorn, Kissing to Find a Match, and SoftSort with our newly introduced ShuffleSoftSort. The evaluation is conducted using 1024 randomly generated RGB colors. For consistency, we use the loss function from [2] for the first three methods, while ShuffleSoftSort employs its adapted version from eq. (2).
The table below compares these methods based on memory requirements for learnable parameters, runtime on an Apple M1 Max, and sorting quality, measured by Distance Preservation Quality (). As demonstrated in [3], is a perceptually driven metric that closely aligns with human visual judgment and strongly correlates with the mean similarity to neighboring elements. Figure 1 shows the images obtained with SoftSort and ShuffleSoftSort.
Method | Memory | Runtime [s] | Quality |
Gumbel-Sinkhorn [11] | 1048576 | 226.8 | 0.913 |
Kissing [4] | 26624 | 114.4 | |
SoftSort [14] | 1024 | 110.7 | 0.698 |
ShuffleSoftSort (ours) | 1024 | 98.0 | 0.892 |
*) invalid permutation |
Further experiments are needed to optimize the parameter settings. However, this table already demonstrates the overall feasibility and effectiveness of the proposed approach, achieving high quality with lowest memory and computational demands. The runtimes are not optimal and only have a relative meaning, as the implementations were not created as optimized GPU applications. ShuffleSoftSort benefits from the fact that its loss function is considerably simpler than that of the other schemes. Notably, the kissing approach exhibits poor convergence due to its simple softmax normalization, often failing to produce valid permutation matrices.
IV Applications
IV-A Grid-based Image Sorting
Viewing large volumes of images is a cognitive challenge, as human perception becomes overwhelmed when too many images are displayed at once. To address this, most applications limit the visible images to around 20. However, sorting images based on the similarity of their visual feature vectors enhances navigation, allowing users to view hundreds of images simultaneously. This method is particularly useful for stock agencies and e-commerce platforms, where efficient browsing is key.
Visual feature vectors, generated through image analysis or deep learning, help organize images by content. Low-level feature vectors, with tens of dimensions, are often more effective for sorting larger image sets, as they enable users to group similar images easily. In this study [3], we used 50-dimensional low-level feature vectors to capture the key visual properties for sorting. Applying grid-based sorting to these vectors results in organized, visually meaningful layouts, as shown below.
IV-B Self-Organizing Gaussians
Another application for large scale permutation learning can be found in the field of 3D scene reconstruction, specifically in 3D Gaussian Splatting (3DGS) [7]. 3DGS has found a lot of attention in recent years, as it solves several challenges in 3D scene reconstruction, such as rendering performance, photorealism, explicitness, or explainability compared to prior methods. On the downside, however, 3DGS suffers from very large storage sizes. This is because 3DGS stores a 3D scene as an unstructured point cloud with millions of data points, each consisting of several parameters like position, scale, orientation, opacity, base color and optionally spherical harmonics. As a result, larger scenes can take up several gigabytes of storage, making them less portable. This has motivated significant research [1] for efficient compression of these datasets. In contrast to other multimedia data, 3DGS data exhibits an interesting property, namely the ambiguity for point ordering. Any reshuffling of Gaussian Splats will lead to exactly the same results, which has been exploited by Self-Organizing Gaussians (SOG) [12]. Instead of storing the data points in vectors, SOG creates a sorted 2D grid for each of the scenes attributes, as shown in Figure 6, increasing spatial correlation of neighboring points. By incorporating a 2D smoothness loss also in the optimization process of the ambiguous splat configurations, further gains can be achieved. The high dimensional maps are then compressed using standard image compression codecs, allowing for up to 40x storage reduction without compromising in rendering quality. The original SOG uses a heuristic non-differentiable sorting, given the large number of data points . However, the proposed method offers gradient-based permutation learning also for such large datasets, requiring only the storage of parameters instead of . This allows for optimally sorting millions of data points without exceeding the memory capacity and enabling end-to-end learning for scene reconstruction.
V Conclusion
In this work, we have introduced ShuffleSoftSort, a novel approach to permutation learning that significantly reduces memory requirements while maintaining high sorting quality. By leveraging an iterative shuffling mechanism, our method improves scalability and enables efficient permutation learning with only parameters, making it well-suited for large-scale optimization tasks. While we have demonstrated alternating sorting in horizontal and vertical directions for grid-based sorting, the method can easily be extended to higher dimensions, with more complex sorting patterns, or alternative shuffling schemes. Compared to existing methods such as Gumbel-Sinkhorn and low-rank approximations, ShuffleSoftSort offers a more memory-efficient solution with only a slight loss of quality. Our results demonstrate that ShuffleSoftSort achieves strong sorting performance while drastically lowering memory demands, making it particularly valuable for applications such as Self-Organizing Gaussians and large-scale data visualization. Future work will explore further optimizations and extensions to multidimensional sorting, reinforcing the potential of differentiable sorting techniques in machine learning and optimization.